### v1

# Coulomb drag driven electron-hole bifluidity in doped graphene

Abstract
Motivated by the notion that a preponderance of Coulomb interactions might lead to hydrodynamics, we carry out an ab initio calculation of the charge carrier transport properties of the electron-hole plasma of doped graphene. We include both the phonon and Coulomb interactions within a momentum and band resolved Boltzmann transport formalism. We find that, under suitable conditions, the strong Coulomb drag effect induces effects like negative conductivity and joint electron-hole hydrodynamics (bifluidity) in the plasma. We also identify the exclusive electron or hole hydrodynamics. We find that there is a strong violation of the Wiedemann-Franz law in the low doped regimes. Our work elucidates the roles of the microscopic scattering mechanisms that drive these hydrodynamic phenomena.
preprint: APS/123-QED
Introduction
The hydrodynamic transport has been a longstanding topic of curiosity in the condensed matter and the broader materials science communities. It is a low dissipation mode of transport, characterized by a large number of scattering events that, however, are (quasi)momentum conserving. In such a non-equilibrium state, a system of particles drifts together with a characteristic global velocity, akin to how water flows through a pipe. This mode of transport lies between the low temperature ballistic and the high temperature diffusive regimes, and has been rather difficult to observe and control in 3D systems, usually due to the inevitable presence of impurity interactions that tend to destroy momentum. 2D materials, on the other hand, allow a way to introduce charge carriers via electrostatic gating, i.e., without the need to introduce dopants. As such, these systems are well-suited for studying hydrodynamics.
Now, graphene [geim2007rise], since its discovery, has been a fertile ground for studying the properties of 2D materials. This is partly because it is easy and inexpensive to obtain, and for its superior charge carrier tunability. It can sustain high charge conduction states [das2011electronic]. And, being a semimetal, it simultaneously hosts both an electron and a hole subsystem. Pertinent to this work is the fact that the charge carrier hydrodynamics in this material has been predicted in several experimental and theoretical studies [crossno2016observation, krishna2017superballistic, sulpizio2019visualizing, muller2009graphene, bistritzer2009hydrodynamic, narozhny2017hydrodynamic, lucas2018hydrodynamics]. From a theoretical point of view, it is interesting to ask what types of scattering processes dictate the charge conductivity of graphene. Interactions of the charge carriers with phonons? With the sample boundary? Interactions between the charge carriers via Coulomb interactions? Since graphene hosts both an electron and a hole gas, it is important to ask whether the alleged hydrodynamic state is only in one or the other subsystem? Or is it of a more exotic kind where both the electron and hole subsystems share a joint global drift velocity, forming what is called as an electron-hole bifluid?
So far, answers to such questions necessarily had to employ semi-empirical transport theories. This is because current publicly available versions of state of the art ab initio Boltzmann electronic transport equation (BTE) solvers [lee2023electron, zhou2021perturbo, protik2022elphbolt, cepellotti2022phoebe] do not provide the functionality to include the interparticle Coulomb interactions in the collision integral. We are aware of one ab initio study on doped silicon for which, however, the software is not publicly available [caruso2016theory]. In that work, the authors include both the electron-electron and electron-plasmon interactions, albeit through a relaxation time approximation (RTA) solution of the BTE. Capturing the hydrodynamic transport physics, however, necessarily requires the full solution of the BTE, or, at least, going beyond the RTA, in order to fully capture the effect of the momentum feedback loop between the two interacting subsystems. To fill the existing methodological and knowledge gaps, we implement the necessary computational tools in the elphbolt [protik2022elphbolt] code to enable an ab initio study of the role of the Coulomb interactions on the charge conduction properties of real materials. Using this newly developed tool, we find that the electron-hole plasma in doped graphene can sustain three types of hydrodynamic states, one of which is the electron-hole bifluid. We find that these hydrodynamic phenomena are driven by the Coulomb interactions between the charge carriers, and that the standard phonon-limited carrier transport theory fails to capture the essential physics, both quantitatively and qualitatively. We further observe a strong Coulomb drag induced negative conductivity, corroborating recent measurements reported in Ref. [ponomarenko2024extreme]. Lastly, we observe a strong violation of the Wiedemann-Franz law close to the bifluid state.
Results and discussion
We operate within the linear response regime where the system is driven slightly out of equilibrium by an external electric field E or a temperature gradient $\mathbf{\nabla}T$. The state of a particle is tagged by a band index $n_{1}$ and wavevector $\mathbf{k}_{1}$, which we write together as $1$. The non-equilibrium distribution of a state is denoted by $f_{1}$, which deviates from the equilibrium, Fermi-Dirac counterpart $f^{0}_{1}$, by a small amount. This small deviation from equilibrium is written in the following form:
$\delta f_{1}=-f^{0}_{1}(1-f^{0}_{1})\beta(\mathbf{\nabla}T\cdot\mathbf{I}_{1}+\mathbf{E}\cdot\mathbf{J}_{1})$ (1)
where $\mathbf{I}_{1}$ and $\mathbf{J}_{1}$ are the response functions to the corresponding applied field and $\beta$ is the inverse temperature energy. BTEs can be written in terms of these response functions, and the solutions then give access to the various transport coefficients, expressions for which have previously been given in Ref. [protik2022elphbolt].
In this work, we focus on capturing the drag effect between electrons and holes, which is the bidirectional flow of momentum between these two systems of charge carriers under non-equilibrium conditions. Since this effect is mediated by the Coulomb interaction, it also goes by the name Coulomb drag in the literature [narozhny2016coulomb]. (Generally, the Coulomb drag is defined as the drag between the charge carrier systems living on macroscopically separated materials. Graphene can be thought of as a special case where the two systems of charge carriers – electron and holes – coexist in the same material sample.) Now, in order to compute this effect within an ab initio framework, we extend the elphbolt code to include a screened Coulomb collision integral. This collision integral has been derived from first principles by Sanborn in Ref. [sanborn1995nonequilibrium] for the case of doped, 3D semiconductors, and here we simply adapt it for our 2D case. In this approximation, the exchange interaction term is ignored. In our notation, the linearized collision integral is given by
$I_{1}^{\text{Coul}}[\mathbf{J}]=$ $\dfrac{4\pi}{\hbar f^{0}_{1}(1-f^{0}_{1})A^{2}}\sum_{23}\sum_{\mathbf{G},\mathbf{G^{\prime}}}\left|\dfrac{V^{\text{2D},\infty}_{\mathbf{q}}}{\epsilon^{\text{TF}}_{\mathbf{q}}(\mathbf{G},\mathbf{G^{\prime}})}\right|^{2}$
$\delta(E_{1}+E_{2}-E_{3}-E_{4})f^{0}_{1}f^{0}_{2}(1-f^{0}_{3})(1-f^{0}_{4})$
$\left(\textbf{J}_{1}+\textbf{J}_{2}-\textbf{J}_{3}-\textbf{J}_{4}\right)\cdot\mathbf{E},$ (2)
where $\hbar$ is the reduced Planck’s constant, $E$ is the energy, $\mathbf{q}\equiv\mathbf{k}_{1}-\mathbf{k}_{3}$, $\hbar\omega\equiv E_{1}-E_{3}$, $A$ is the area of the crystal, and $V^{\text{2D},\infty}_{\mathbf{q}}$ is the 2D bare Coulomb interaction[saraga2004coulomb] screened by a high frequency dielectric. Note that the wave vector of state 4 is determined by those of the remaining three. We choose the value of $\epsilon^{\infty}$ to be that of an hBN substrate (4.14.1)[lin2013surface] to model the experimental conditions of the work in Ref. [ponomarenko2024extreme]. $\epsilon^{\text{TF}}_{\mathbf{q}}$ is the Thomas-Fermi dielectric, given by
$\epsilon_{\mathbf{q}}^{\text{TF}}(\mathbf{G},\mathbf{G}^{\prime})=\delta_{\mathbf{G},\mathbf{G}^{\prime}}+\dfrac{e^{2}\beta}{2\varepsilon_{0}A}\sum_{1}\dfrac{f^{0}_{1}(1-f^{0}_{1})}{\sqrt{|\mathbf{q}+\mathbf{G}||\mathbf{q}+\mathbf{G}^{\prime}|}},$ (3)
where $e$ is the magnitude of the electronic charge and $\varepsilon_{0}$ is the permittivity of free space. The corresponding expression for the temperature gradient field case is found through the replacement ($\mathbf{J}\rightarrow\mathbf{I}$, $\mathbf{E}\rightarrow\mathbf{\nabla}T$). Going forward, we ignore all local field effects, i.e., set $\mathbf{G}=\mathbf{G}^{\prime}$.
The Coulomb collision integral is added to the electron-phonon counterpart [protik2022elphbolt] assuming Matthiesen’s rule, i.e., that the scattering channels are independent. The final charge carrier BTE is fully solved to self-consistency using an iterative approach. That is, our approach is beyond any type of RTA. Doing so is crucial in order to accurately capture any hydrodynamics. Details of our ab initio BTE solution method have been given before in Ref. [protik2022elphbolt] and are not reproduced here. But, briefly, this approach relies on the density functional + perturbation theory as implemented in the Quantum Espresso code [giannozzi2009quantum, giannozzi2017advanced], maximally localized Wannier functions [pizzi2020wannier90], and the real-space, Wannier representations of the electron-phonon interaction matrix elements obtained by interfacing with the EPW code [giustino2007electron, ponce2016epw]. Details of the employed computational methods are given in the Supplementary Information (SI) [suppl].
Since we compare to the measurements in Ref. [ponomarenko2024extreme], we include a boundary scattering term in all our calculations using a simple, phenomenological scattering rates expression: $|\mathbf{v}_{1}|/L$, where $\mathbf{v}_{1}$ is the charge carrier group velocity and $L=20$ $\mu\text{m}$ characterizes the sample size. This expression is added to the rest of the RTA scattering rates using, again, Matthiesen’s rule. We ignore any in-scattering corrections from this scattering channel.
In this work, we choose the majority carrier concentrations to range between $1.6\times 10^{11}$ and $1.7\times 10^{12}$ cm-2. Getting any closer to the charge neutrality point is not possible at the moment due to the extreme computational complexity; see SI [suppl]. In any case, the application of the semiclassical BTE is invalid near the charge the neutrality point [castro2009electronic].
Figure 1: Resistivity vs. charge carrier concentration at various temperatures. Positive (negative) sector of the horizontal axis corresponds to majority electron (hole) concentration. Solid green line gives the 300 K measurements from Ref. [ponomarenko2024extreme] (data obtained from the primary author via private communication and converted to resistivity units). Calculated values are presented in disks for three temperatures: 200 K (blue), 300 K (green), and 400 K (red). Solid (empty) symbols correspond to the theory that includes (excludes) the Coulomb interactions.
In Fig. 1, we present our calculated room temperature resistivity as a function of the difference of the electron and hole concentrations, $n_{e}-n_{h}$. We include a comparison to recent $300$ K measurements from Ref. [ponomarenko2024extreme] (green line). We find that without the inclusion of the Coulomb interactions (empty disks), the calculated resistivity is in poor agreement with the measurements (green line vs. green empty disks) over the entire range of concentrations considered. The theory including the Coulomb interactions gives quantitatively good agreement (green line vs. green solid disks) in the moderate to high concentration range. For low concentrations, however, the Coulomb enabled theory overestimates the resistivity. This could be due to a couple of reasons. First, the lattice constant used in our theoretical modeling might not match the one of the sample in Ref. [ponomarenko2024extreme] since the graphene flake there is sandwiched between hBN substrates. We take into account the role of hBN only through an extra screening factor, as mentioned earlier. Secondly, our Coulomb collision integral includes only the direct interaction term. Our theoretical predictions might be improved through the inclusion of the exchange term, since the direct and exchange terms have opposite signs. We leave this for a future work. Now, the transport coefficients presented in Fig. 1 being band summed quantities, hide certain crucial failures of the Coulomb-free theory. And in order to observe these failures of the phonon limited transport to capture important microscopic physics, we next look at the valence and conduction band contributions separately.
Figure 2: $200$ K spectral conductivity from the valence and conduction bands for hole doped cases. Thick (thin) lines are for a majority hole concentration of $1.3\times 10^{12}$ ($2\times 10^{11}$) cm-2. Solid (dashed) lines denote the theory including (excluding) the Coulomb interactions. The Dirac point is set at zero energy. The contribution from the conduction band (positive energy) is displayed with a factor $10$ enlargement.
In Fig. 2, we give the spectral conductivities at $200$ K for a high (thick lines) and a low (thin lines) hole doped case. The Dirac point is placed at the zero of the horizontal axis. Note that the conduction band contribution is plotted with a scaling factor of $10\times$. For the Coulomb-free case (dashed lines), both the valence and the conduction bands give positive contributions to the conductivity, as one would expect from a classical Drude picture. In stark contrast, the activation of the Coulomb interactions (solid lines) leads to a significantly lower positive contribution from the valence band, along with a negative contribution from the conduction band. A completely analogous effect happens for the electron-doped case, where the inclusion of the Coulomb interactions causes the holes to give a negative contribution to the conductivity; see SI [suppl]. There we also show that this effect persists even at room temperature. What we described so far is the evidence for an extreme Coulomb drag effect between the electron and hole systems. The essential physics we observe here is that the majority carriers transfer momentum to the minority ones, causing them to develop a current in a direction opposite to what is dictated by the coupling to the external electric field. On the other hand, the minority carriers also drag the majority carriers in the opposite direction, causing them to give a lower, albeit still positive, contribution. This extreme Coulomb drag effect becomes more prominent the closer one gets to the charge neutrality point. Note that this phenomenon has been experimentally observed in Ref. [ponomarenko2024extreme]. It is now interesting to inquire whether the mode of this charge transport is hydrodynamic. This question is motivated by the prevailing wisdom that the Coulomb scatterings between the charge carriers contain a large proportion of the momentum conserving (Normal) type processes. And in systems where Normal processes dominate over momentum destroying (Umklapp or impurity) ones, hydrodynamic behavior manifests. In such a case, the non-equilibrium distribution function takes the form of a shifted Fermi distribution, the shift amount being characterized by a global drift velocity. We investigate this next.
(a) T = $200$ K, $n_{h}-n_{e}$ = $2.1\times 10^{11}\text{ cm}^{-2}$
(b) T = $200$ K, $n_{h}-n_{e}$ = $7.4\times 10^{11}\text{ cm}^{-2}$
(c) T = $200$ K, $n_{h}-n_{e}$ = $1.7\times 10^{12}\text{ cm}^{-2}$
(d) T = $200$ K
[…]
Figure 4: Lorentz number vs. carrier concentration. Blue, green, and red symbols are for $200$, $300$, and $400$ K, respectively. Solid (broken) lines correspond to the theory including (excluding) the Coulomb interaction. The dashed horizontal line gives the universal Lorentz number ($2.44\times 10^{-8}\text{ V}^{2}\text{ K}^{-2}$) according to the Wiedemann-Franz law.
Finally, in Fig. 4, we plot the Lorentz number as a function of $n_{e}-n_{h}$, for three temperature points. The horizontal dashed black line gives the universal Lorentz number. The prevailing wisdom is that very close to the charge neutrality point in graphene at low temperatures, a violation of the Wiedemann-Franz (WF) law occurs due to the domination of the electron-hole interactions [crossno2016observation]. Here we find that when the Coulomb interactions are included (solid blue, green, and red lines), there is a strong violation of the WF law even in the electron and hole doped regimes. However, this violation increases as we approach the charge neutrality. This, however, does not correlate with the bifluidity, which we found to be a dominant effect at low temperatures, while here, the WF law violation increases at high temperatures. The theory that ignores the Coulomb interactions (dashed blue, green, and red lines), also shows a violation of the WF law, though much weaker than what is predicted by the better theory.
Summary and outlook
In this work we implement an ab initio method to include the Coulomb interactions among the charge carriers within the elphbolt Boltzmann transport codes suite. We use this new tool to compute the charge carrier transport properties of doped graphene. We find that the Coulomb interaction enabled theory predicts three hydrodynamic transport regimes: (1) exclusive hole hydrodynamics in the high hole doped case, (2) exclusive electron hydrodynamics in the high electron doped case, and (3) coupled electron-hole bilfluidity in the weak hole or electron doped cases at low temperature. Our results corroborate the recent observations of negative conductivity in the extreme Coulomb drag regime [ponomarenko2024extreme]. We also find a strong violation of the Wiedemann-Franz law that is driven by the Coulomb interactions between electrons and holes. None of the above mentioned phenomena is captured by the theory that includes the electron-phonon interactions, but ignores Coulomb.
Although the ab initio methodology described in this work is able to predict the strong Coulomb drag driven negative conductivity and the electron-hole bifluidity, we see several avenues for improvement and extensions that we wish to touch upon in future works. We briefly discuss them below.
  • Exchange term: As mentioned earlier, our collision integral contains only the direct term. Since the direct and exchange terms come with opposite signs, we think that the inclusion of the latter will reduce calculated resistivity values at low doping levels, and result in theoretical predictions that are in better agreement with measurements.
  • Dynamical screening: In this work, we employ the Thomas-Fermi screening model. One can, in fact, use a dynamical screening model within the random phase approximation (RPA) [sanborn1995nonequilibrium]. The RPA includes the effect of the plasmons in the Coulomb kernel and, in principle, captures the screening better than the static Thomas-Fermi theory.
  • Electron-phonon drag: In this work, we assume that the phonon system can be taken to remain in equilibrium. We do this for two reasons. Firstly, we notice that the transport physics in the doping and temperature ranges considered here is entirely dictated by the Coulomb interactions, as is evidenced from the failure of the theory that ignores these. Secondly, computing the phonon thermal conductivity in graphene is a difficult and contentious matter. It has been argued that including the extremely computationally demanding four-phonon interactions is necessary to make decent lattice thermal conductivity predictions. That said, there might as well be other systems where the extremely expensive four-phonon interactions do not have to be considered. In such system, it might be possible to carry out a coupled electron-phonon BTEs calculation including the Coulomb interactions. We note that in Ref. [quan2025coupled], the authors predict a joint electron-phonon hydrodynamics in doped MoS2. An interesting research question would then be whether there exist materials that, under suitable conditions, exhibit both the electron-hole and electron-phonon hydrodynamics.
Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Emmy Noether research grant (Grant No. 534386252). This work used computational resources provided by Noctua2 at the Paderborn Center for Parallel Computing (PC2), JUWELS at the Jülich Supercomputing Centre, and the CMS cluster of Humboldt-Universität zu Berlin. We thank Dr. Leonid A. Ponomarenko for kindly providing the resistivity measurement from Ref. [ponomarenko2024extreme].
Data and code availability
The theory results supporting the findings of this work will be made available at the time of publication. The elphbolt code used to carry out the calculations is free/libre and open source software.